one share of A stock is worth 3 1/2 shares of B stock. If the total value of the stock is $9000, how much was invested in each company?
Answer by kiru_khandelwal(79)

Let the number of A stock be a
and the number of B stock be b
Now one share of A stock is worth 31/2 shares of B stock
=> a = 3 1/2 b
=> a = 7/2b
Now
a + b = 9000
=> 7/2b + b = 9000
=> 9/2b = 9000
=> b = 2000
=> a = 7/2*2000 = 7000
So in stock A $7000 is invested and in stock B $2000 is invested